OpenStudy (nick88888888):

Find the equation in slope-intercept form of the line through points (-2, 5) and (6, -15). How would i do this? Do i use y = mx+b? or what im not sure how to do it

OpenStudy (16shuston):

ya use y=mx+b

OpenStudy (16shuston):

do you know how to figure out the slope?

OpenStudy (nick88888888):

No ;-;

OpenStudy (16shuston):

okay so it will be change in y over change in x

OpenStudy (nick88888888):

I mean i was asking cause i was trying to search how to do it and some would say y = mx+b and then some would be solving it like this (-2,5) and (6, -15) First find slope. (m) (y2-y1)/(x2-x1) m=(-15-5)/(6-2)

OpenStudy (16shuston):

ya you are doing it right so far

OpenStudy (16shuston):

so then when you find m you use a set o your points to find b.

OpenStudy (16shuston):

do you want me to show you?

OpenStudy (nick88888888):

wait so like this? (-2,5) and (6, -15) First find slope. (m) (y2-y1)/(x2-x1) m=(-15-5)/(6-2) m=-20/4 is the slope now replace x and y terms (y = mx+b) 5=-20/4*2+b? is that correct so far?

OpenStudy (triciaal):

x2 -x1 = 6 - (-2) so it becomes 6 + 2

OpenStudy (nick88888888):

oh.. so (-2,5) and (6, -15) First find slope. (m) (y2-y1)/(x2-x1) x2 -x1 = 6 - (-2) so it becomes 6 + 2 m=(-15-5)/(6+2) m=-20/8 is the slope now replace x and y terms (y = mx+b) 5=-20/8*2+b?

OpenStudy (triciaal):

so the slope is -20/8 = -5/2 using one of the given points, (x1, y1) and (x,y) y-y1 = -5/2(x-x1)

OpenStudy (nick88888888):

This is how a certain site said to do it :/ (-2,5) and (6, -15) First find slope. (m) (y2-y1)/(x2-x1) x2 -x1 = 6 - (-2) so it becomes 6 + 2 m=(-15-5)/(6+2) m=-20/8 is the slope now replace x and y terms (y = mx+b) 5=-20/8*2+b? 5=-40/8+b 5+40/8=b (40+40)/8=b 80/8=b the y intercept Thus equation being y=-20x/8+80/8

OpenStudy (triciaal):

@Nick88888888 you don't need a site to tell you to always reduce to the lowest term

OpenStudy (butterflydreamer):

hmm... not quite. First you want to find the slope (m) So, use the formula:  Then you want to use the formula: y = mx + b (this is your point-intercept formula) Use (-2, 5) where (x,y) therefore x = -3, y = 5 and we know m(slope) = -20/8 Therefore: \[y=mx + b \rightarrow 5= -20/8 \times-2 +b \] From here, multiply everything by 8 to get rid of the fraction: 5*8 = -20(-2) + 8b 40 = 40 + 8b Subtract 40 from both sides, 40 - 40 = 40 + 8b - 40 0 = 8b Find the value of b by dividing both sides by 8 and we get b = 0 Then plug this into your formula: y= mx + b , where we now know the value of m and b :)